# May 2, 2024

### v1

@@ Abstract
Integrable deformations of a class of Rikitake dynamical systems are constructed by deforming their underlying Lie-Poisson Hamiltonian structures as linearizations of Poisson–Lie structures on certain (dual) Lie groups. By taking into account that there exists a one-to one correspondence between Poisson–Lie groups and Lie bialgebra structures, a number of deformed Poisson coalgebras can be obtained, which allow the construction of integrable deformations of coupled Rikitake systems. Moreover, the integrals of the motion for coupled systems can be explicitly obtained by means of the deformed coproduct map. The same procedure can be also applied when the initial system is bi-Hamiltonian with respect to two different Lie-Poisson algebras. In this case, to preserve a bi-Hamiltonian structure under deformation, a common Lie bialgebra structure for the two Lie-Poisson structures has to be found. Coupled dynamical systems arising from this bi-Hamiltonian deformation scheme are also presented, and the use of collective ‘cluster variables’, turns out to be enlightening in order to analyse their dynamical behaviour. As a general feature, the approach here presented provides a novel connection between Lie bialgebras and integrable dynamical systems.
@@ 1 Introduction
$\left\{f,g\right\}(\xi)=\left\langle\xi,\left[df,dg\right]\right\rangle,\qquad\xi\in\mathfrak{g}^{*},\quad f,g\in C^{\infty}(\mathfrak{g}^{*}),$
where $\langle$,$\rangle$ , ⟩ is the pairing between $\mathfrak{g}$ and $\mathfrak{g}^{*}$. This Lie–Poisson bracket is given by a set of fundamental Poisson brackets which are just the Poisson analogues of the Lie brackets defining $\mathfrak{g}$.
Let us now consider a generic $N$-dimensional (ND) dynamical system defined by the Lie-Poisson structure associated to a given finite-dimensional Lie algebra $\mathfrak{g}$ together with a Hamiltonian function $\cal{H}\in C^{\infty}(\mathfrak{g}^{*})$. The Hamilton equations of motion of the system will be given by the system of ODEs
@@
where $x_{i}$ are coordinate functions on $\mathfrak{g}^{*}$. By construction, the set of Casimir functions ${\cal C}_{i}$ of the Lie-Poisson structure provides a set of integrals of the motion in involution for such a system, and level sets of the Casimir functions define the symplectic leaves of the Lie-Poisson structure.
The problem of finding integrable deformations of such a system admits many solutions. A first obvious family of integrable deformations can be obtained just by modifying the Hamiltonian $\cal{H}$ and by preserving the Lie-Poisson algebra and its Casimir functions ${\cal C}_{i}$. Indeed, more interesting integrable deformations will be the ones obtained by considering nontrivial deformations of the underlying Lie-Poisson structure together with the associated deformed versions of the Casimir functions.
@@
The aim of this paper is to show that, among this infinite zoo of possible integrable deformations, there exists a distinguished and very restricted subclass of deformations that preserves the additional Poisson coalgebra symmetry that exists for any Lie-Poisson dynamical system, which was introduced in [2] and is based on the existence of a coproduct map that endows the Lie-Poisson algebra with a Poisson coalgebra structure. In this framework, the integrable deformation arises as a consequence of the existencie of a compatible deformed version of the coproduct map, which is just the group law of a dual non-abelian Poisson-Lie group. Moreover, the essential property of such Poisson coalgebra deformations is the fact that the deformed coproduct map provides coupled pairs of deformations of the initial system in a systematic and constructive way, which was introduced in [2, 3, 4, 5], latter applied to 3D Lotka-Volterra systems in [6] and further generalised to some bi-Hamiltonian systems in [7, 8]. In this framework, the deformed coproduct map of the deformed Casimir functions will give rise to the integrals of the motion of the deformed system, thus providing its complete integrability (in the Liouville sense) structure.
@@
The paper is organized as follows. In the next Section the class of integrable Rikitake-type systems is reviewed, together with the Lie-Poisson Hamiltonian structures that underly their integrability. Specifically, we will select the three types of integrable Rikitake systems that will be considered in the paper (called A, B and AB). In Section 3, the interpretation of Poisson-Lie groups as (coalgebra) deformations of Lie-Poisson algebras is presented, and the construction will be illustrated through the specific case of the Lie-Poisson (1+1) Poincaré algebra, which will be the relevant one in the rest of the paper. In Section 4, the Poincaré Poisson coalgebra symmetry of an integrable Rikitake system of type A will be presented. Furthermore, we prove that bi-Hamiltonian deformations with Poisson coalgebra symmetry does not exist for the Rikitake system of type B. Finally, a Poisson-Lie group structure suitable for the construction of a bi-Hamiltonian deformation for the Rikitake AB system will be presented. In Section 5, we compute the integrable deformations of the Rikitake system arising from the Poisson-Lie structures constructed in Section 4. In Section 6, the coupled systems arising from the Poisson-Lie symmetry for the Rikitake AB system will be explicitly obtained. Finally, a section including some comments and open problems closes the paper.
@@ 2 Lie-Poisson algebras for Rikitake systems
where $\mu,\beta_{1},\beta_{2}$ and $\alpha$ are real constants (the original Rikitake system [12] is the one with $\beta_{2}=0$). The integrability properties of the system (2.1) have been thouroughly studied [14, 15, 16] (see also [17, 18, 19, 20, 21] and references therein). Moreover, (integrable) deformations of the Rikitake system have attracted a considerable attention (see [22, 23, 24, 25]).
There exist two distinguished families of integrable cases for this dynamical system (A and B), where $\mu=0$ arises always as a common integrability condition. In the sequel we describe both integrable families together with their associated Lie-Poisson algebras, and we shortly review the ways in which integrable deformations for them can be constructed.
@@ 2.1 Case A
$\begin{array}[]{lll}\dot{x}=y\,z,&\quad\dot{y}=x\,z,&\quad\dot{z}=\alpha-xy.\end{array}$ (2.2)
@@
defines a map from the whole Poisson manifold to a given simplectic leaf, and allows us to write the Hamiltonian (2.3) in terms of the canonical variables $q,p$, namely
@@ 2.2 Case B
$\begin{array}[]{lll}\dot{x}=y(z+\beta),&\quad\dot{y}=x(z-\beta),&\quad\dot{z}=-xy.\end{array}$ (2.10)
@@
$\{x,y\}_{0}=2\,\beta\,I\,\qquad\{y,z\}_{0}=x,\qquad\{z,x\}_{0}=-y,\qquad\{I,\cdot\}_{0}=0,$ (2.11)
@@
{.,.}$\quad\{.,.\}_{\lambda}=(1-\lambda)\{.,.\}_{0}+\lambda\{.,.\}_{1},\quad\mbox{ with }\lambda\in\mathbb{R}$ (2.17)
whose explicit brackets are given by
@@
Therefore, the approach introduced in [8] (see also [26]) shows that for each common 1-cocycle for the 4D Lie algebras (2.11) and (2.14) we could, in principle, construct an integrable bi-Hamiltonian deformation of the system. However, in this particular case, we will show that such a common cocycle does not exist. Nevertheless, the usual integrable deformation procedure that we apply to the case A can be also used in this case, obtaining a different deformation for each of the Lie–Poisson structures.
@@ 2.3 Case AB
Nevertheless, when exploring the possible common cocycles for the Poisson pencil (2.18) we realize that the case $\beta=0$ do admit such a bi-Hamiltonian structure with a common cocycle (note that this is also Case A with $\alpha=0$). This system, namely
@@
Then, by following [8] we will construct a bi-Hamiltonian deformation of (2.19) constructed through the Poisson pencil
@@
Also, the integrable coupled systems arising from the coalgebra approach will be explicitly presented. The Casimir operator for (2.25) has the following expression:
@@
$x=\sqrt{4\,k+2\,\lambda\,p^{2}}\,\sinh(\sqrt{1-2\lambda}\,q),\qquad y=\sqrt{\dfrac{4\,k+2\,\lambda\,p^{2}}{1-2\,\lambda}}\,\cosh(\sqrt{1-2\lambda}\,q),\qquad z=p$ (2.27)
@@
For the second one, corresponding to setting $\lambda=1$ in (2.27), the Hamiltonian (2.23) in canonical coordinates is given by
@@ 3 Lie bialgebras
Lie bialgebras are the tangent counterpart of Poisson-Lie groups [11] in an similar way Lie algebras are the tangent structures associated to Lie groups. In our case, the extra Poisson structure on the Lie group $G$ induces the cocommutator map $\delta$ on $T_{e}G\simeq\mathfrak{g}$. This implies that, for any dynamical system that admits a Hamiltonian formulation in terms of a linear Poisson structure (Lie-Poisson structure), each possible bialgebra structure for the associated Lie algebra gives rise to a deformed Hamiltonian system. The remarkable properties of deformed systems so constructed stems from the fact that the algebra of smooth functions $\mathcal{C}^{\infty}(G)$ on the Poisson-Lie group $G$ is automatically endowed with a Poisson-Hopf algebra structure, characterized by the coproduct map $\Delta:\mathcal{C}^{\infty}(G)\to\mathcal{C}^{\infty}(G)\otimes\mathcal{C}^{\infty}(G)$. This implies that deformed systems constructed in this way are very rigid, compared to arbitrary deformations of the Poisson structure and/or the Hamiltonian. However, this Poisson-Hopf structure automatically guarantees that canonically coupled systems can be constructed by means of the coproduct map $\Delta$. For a detailed explanation of all these results, see [7, 8] and references therein.
3.1 Lie bialgebra structures for the $(1+1)$ Poincaré Lie algebra
Let us consider the $(1+1)$-dimensional Poincaré Lie algebra
$\begin{array}[]{lll}[X,Y]=0,&[X,Z]=Y,&[Y,Z]=X.\end{array}$ (3.5)
which is relevant for Case A, Case B (when centrally extended) and Case AB. The full classification of nonisomorphic Lie bialgebra structures for this algebra is given in [28] and leads to seven non-isomorphic families of cocommutator maps. Out of them, let us consider the following three which give rise to different kinds of dual Lie algebras: the trivial one whose dual Lie algebra is Abelian, and two non-trivial cocommutators, one with nilpotent and one with solvable dual Lie algebras.
@@ 3.1 Lie bialgebra structures for the $(1+1)$ Poincaré Lie algebra
    with dual Lie algebra isomorphic to the ‘book’  Lie algebra
@@
    with dual Lie algebra isomorphic to the Heisenberg Lie algebra, namely
@@
This means that the three dual Lie bialgebras $(\mathfrak{g}^{\ast},\delta^{\ast})$ will have the same dual cocommutator map $\delta^{*}$ arising from the Poincaré Lie algebra
@@
4 Poisson-Lie deformations of the Poincaré algebra
In this Section we study the Poisson-Hopf algebras relevant to the different deformations of the Rikitake systems that we are considering. In particular, for Case A we explicitly compute both of the non-trivial Poisson-Hopf algebras associated to the two non-trivial $(1+1)$-dimensional Poincaré Lie bialgebra introduced below and show explicitly that in the limit $\eta\to 0$ we recover the Poisson version of (3.5). Afterwards, we study the possible bi-Hamiltonian deformations: we prove that Case B does not admit a bi-Hamiltonian deformation by showing that the Poisson pencil (2.18) does not admit a common cocycle, and finally, we explicitly construct a bi-Hamiltonian deformation for Case AB.
@@ 4.1.1 The ‘book’ group deformation
By using (4.1) we can parametrize a general element of the group $G^{*}$ in terms of the $(x,y,z)$-coordinates, namely
$\mathcal{G}_{1}=e^{z\,\varrho(\hat{z})}\cdot e^{y\,\varrho(\hat{y})}\cdot e^{x\,\varrho(\hat{x})}=\begin{pmatrix}e^{-\eta\,z}&0&\eta\,xe^{-\eta z}\\ 0&e^{-\eta\,z}&\eta\,ye^{-\eta z}\\ 0&0&1\end{pmatrix},$ (4.2)
@@
From (4.3) the coproducts for the $(x,y,z)$-coordinates read
$\begin{array}[]{l}\Delta_{\eta}(x)=x\otimes e^{\eta\,z}+1\otimes x=x_{1}e^{\eta\,z_{2}}+x_{2}\\ \Delta_{\eta}(y)=y\otimes e^{\eta\,z}+1\otimes y=y_{1}e^{\eta\,z_{2}}+y_{2}\\ \Delta_{\eta}(z)=z\otimes 1+1\otimes z=z_{1}+z_{2}\,,\end{array}$ (4.4)
where we have denoted the coordinates of $\mathcal{G}_{1}$ with the tensor space at the left side of the tensor product, and the ones of $\mathcal{G}_{2}$ at the right one.
As the deformed Poisson bracket is quadratic in terms of the group entries, the application of the compatibility conditions stated by the fact that the deformed coproduct relations (4.4) have to provide a Poisson map leads to the most generic PL structure on the book group, which was explicitly obtained in [27].
By imposing that the linearisation of such a generic PL structure leads to a PL Poincaré algebra, a straightforward computation leads to the Poisson bracket
@@
4.1.2 The Heisenberg group deformation

### v2

# June 18, 2024

@@ Abstract
Integrable deformations of a class of Rikitake dynamical systems are constructed by deforming their underlying Lie-Poisson Hamiltonian structures, which are considered linearizations of Poisson–Lie structures on certain (dual) Lie groups. By taking into account that there exists a one-to one correspondence between Poisson–Lie groups and Lie bialgebra structures, a number of deformed Poisson coalgebras can be obtained, which allow the construction of integrable deformations of coupled Rikitake systems. Moreover, the integrals of the motion for these coupled systems can be explicitly obtained by means of the deformed coproduct map. The same procedure can be also applied when the initial system is bi-Hamiltonian with respect to two different Lie-Poisson algebras. In this case, to preserve a bi-Hamiltonian structure under deformation, a common Lie bialgebra structure for the two Lie-Poisson structures has to be found. Coupled dynamical systems arising from this bi-Hamiltonian deformation scheme are also presented, and the use of collective ‘cluster variables’, turns out to be enlightening in order to analyse their dynamical behaviour. As a general feature, the approach here presented provides a novel connection between Lie bialgebras and integrable dynamical systems.
@@ 1 Introduction
$\left\{f,g\right\}(\xi)=\left\langle\xi,\left[df,dg\right]\right\rangle,\qquad\xi\in\mathfrak{g}^{*},\quad f,g\in C^{\infty}(\mathfrak{g}^{*}),$ (1.1)
[…]
Let us now consider a generic $N$-dimensional (ND) dynamical system defined by the Lie-Poisson structure associated to a given finite-dimensional Lie algebra $\mathfrak{g}$, together with a Hamiltonian function $\cal{H}\in C^{\infty}(\mathfrak{g}^{*})$. The Hamilton equations of motion of the system will be given by the system of ODEs
@@
where $x_{i}$ are coordinate functions on $\mathfrak{g}^{*}$. By construction, the set of Casimir functions ${\cal C}_{i}$ of the Lie-Poisson structure provides a set of integrals of the motion in involution for such a system, and level sets of the Casimir functions define the symplectic leaves of the Lie-Poisson structure (1.1).
The problem of finding integrable deformations of the system admits many solutions. A first obvious family of integrable deformations can be obtained by modifying the Hamiltonian $\cal{H}$ and by preserving the Lie-Poisson algebra (1.1) and its Casimir functions ${\cal C}_{i}$. Indeed, more interesting integrable deformations will be the ones obtained by considering nontrivial deformations of the underlying Lie-Poisson structure together with the associated deformed versions of the Casimir functions.
@@
The aim of this paper is to show that, among this infinite zoo of possible integrable deformations, there exists a distinguished and very restricted subclass of deformations that preserves the additional Poisson coalgebra symmetry that exists for any Lie-Poisson dynamical system, which was introduced in [2] and is based on the existence of a coproduct map that endows the Lie-Poisson algebra with a Poisson coalgebra structure. In this framework, the integrable deformation arises as a consequence of the existence of a compatible deformed version of the coproduct map, which is just the group law of a dual non-abelian Poisson-Lie group. Moreover, the essential property of such Poisson coalgebra deformations is the fact that the deformed coproduct map provides coupled pairs of deformations of the initial system in a systematic and constructive way, which was introduced in [2, 3, 4, 5], latter applied to 3D Lotka-Volterra systems in [6] and further generalised to some bi-Hamiltonian systems in [7, 8]. In this framework, the deformed coproduct map of the deformed Casimir functions will give rise to the integrals of the motion of the deformed system, thus providing its complete integrability structure in the Liouville sense.
@@
The paper is organized as follows. In the next Section the class of integrable Rikitake-type systems is reviewed, together with the Lie-Poisson Hamiltonian structures that underly their integrability. Specifically, we will select the three types of integrable Rikitake systems that will be considered in the paper (called A, B and AB). In Section 3, the interpretation of Poisson-Lie groups as (coalgebra) deformations of Lie-Poisson algebras is presented, and the construction will be illustrated through the specific case of the Lie-Poisson (1+1) Poincaré algebra, which will be the relevant one in the rest of the paper. In Section 4, the Poincaré Poisson coalgebra symmetry of an integrable Rikitake system of type A will be presented. Furthermore, we prove that bi-Hamiltonian deformations with Poisson coalgebra symmetry does not exist for the Rikitake system of type B. Finally, a Poisson-Lie group structure suitable for the construction of a bi-Hamiltonian deformation for the Rikitake AB system will be presented. In Section 5, we compute the integrable deformations of the Rikitake system arising from the Poisson-Lie structures constructed in Section 4. In Section 6, the coupled systems coming from the Poisson-Lie symmetry for the Rikitake AB system will be explicitly obtained. Finally, a section including some comments and open problems closes the paper.
@@ 2 Lie-Poisson algebras for Rikitake systems
where $\mu,\beta_{1},\beta_{2}$ and $\alpha$ are real parameters (the original Rikitake system [12] is the one with $\beta_{2}=0$). The integrability properties of the system (2.1) have been thouroughly studied [14, 15, 16] (see also [17, 18, 19, 20, 21] and references therein). Moreover, (integrable) deformations of the Rikitake system have attracted a considerable attention (see [22, 23, 24, 25]).
There exist two distinguished families of integrable cases for this dynamical system (A and B), where $\mu=0$ arises always as a common integrability condition. In the sequel we describe both integrable families together with their associated Lie-Poisson algebras, and we sketch the ways in which integrable deformations for them can be constructed.
@@ 2.1 Case A
$\begin{array}[]{lll}\dot{x}=y\,z,&\quad\dot{y}=x\,z,&\quad\dot{z}=\alpha-xy,\end{array}$ (2.2)
@@
defines a map from the whole Poisson manifold to a given simplectic leaf, and allows us to write the Hamiltonian (2.3) in terms of the canonical variables $(q,p)$, namely
@@ 2.2 Case B
$\{x,y\}_{0}=2\,\beta\,I,\qquad\{y,z\}_{0}=x,\qquad\{z,x\}_{0}=-y,\qquad\{I,\cdot\}_{0}=0,$ (2.11)
@@
{.,.}$\quad\{.,.\}_{\lambda}=(1-\lambda)\{.,.\}_{0}+\lambda\{.,.\}_{1},\quad\mbox{ with }\lambda\in\mathbb{R},$ (2.17)
whose explicit Poisson brackets are given by
@@
Therefore, the approach introduced in [8] (see also [26]) shows that for each common 1-cocycle for the 4D Lie algebras (2.11) and (2.14) we could, in principle, construct an integrable bi-Hamiltonian deformation of the system. However, in this particular case, we will show in Section 4.2 that such a common cocycle does not exist. Nevertheless, the usual integrable deformation procedure that we apply to the case A can be also used in this case, thus obtaining a different deformation for each of the Lie–Poisson structures.
@@ 2.3 Case AB
Nevertheless, when exploring the possible common cocycles for the Poisson pencil (2.18) we realize that the case $\beta=0$ does admit such a bi-Hamiltonian structure with a common cocycle (note that this is also Case A with $\alpha=0$). This system, namely
@@
Then, by following [8] we will construct a bi-Hamiltonian deformation of (2.19) by means of the Poisson pencil
@@
Also, the integrable coupled systems arising from the coalgebra approach will be explicitly presented. Note that the Casimir function for (2.25) has the following expression:
@@
$x=\sqrt{4\,k+2\,\lambda\,p^{2}}\,\sinh(\sqrt{1-2\lambda}\,q),\qquad y=\sqrt{\dfrac{4\,k+2\,\lambda\,p^{2}}{1-2\,\lambda}}\,\cosh(\sqrt{1-2\lambda}\,q),\qquad z=p.$ (2.27)
@@
For the second one, corresponding to setting $\lambda=1$ in (2.27), the symplectic realization of the Hamiltonian (2.23) is given by
@@ 3 Lie bialgebras: the $(1+1)$ Poincaré case
Lie bialgebras are the tangent counterpart of Poisson-Lie groups [11] in an similar way that Lie algebras are the tangent structures associated to Lie groups. In our case, the extra Poisson structure on the Lie group $G$ induces the cocommutator map $\delta$ on $T_{e}G\simeq\mathfrak{g}$. This implies that, for any dynamical system that admits a Hamiltonian formulation in terms of a linear Poisson structure (Lie-Poisson structure), each possible bialgebra structure for the associated Lie algebra gives rise to a deformed Hamiltonian system. The remarkable properties of deformed systems so constructed stem from the fact that the algebra of smooth functions $\mathcal{C}^{\infty}(G)$ on the Poisson-Lie group $G$ is automatically endowed with a Poisson-Hopf algebra structure, characterized by the coproduct map $\Delta:\mathcal{C}^{\infty}(G)\to\mathcal{C}^{\infty}(G)\otimes\mathcal{C}^{\infty}(G)$. This implies that deformed systems constructed in this way are very rigid, compared to arbitrary deformations of the Poisson structure and/or the Hamiltonian. However, this Poisson-Hopf structure automatically guarantees that canonically coupled systems can be constructed by means of the coproduct map $\Delta$. For a detailed explanation of all these results, see [7, 8] and references therein.
Let us now consider the $(1+1)$-dimensional Poincaré Lie algebra
$\begin{array}[]{lll}[X,Y]=0,&[X,Z]=Y,&[Y,Z]=X,\end{array}$ (3.5)
which is relevant for Case A, Case B (when the central extension is considered) and Case AB. The full classification of nonisomorphic Lie bialgebra structures for this algebra is given in [28] and leads to seven non-isomorphic families of cocommutator maps. Out of them, we will consider the trivial one whose dual Lie algebra is Abelian, and two non-trivial cocommutators, one with nilpotent and one with solvable dual Lie algebras. Explicitly:
@@
    with dual Lie algebra isomorphic to the ‘book’  Lie algebra [27]
@@
    with dual Lie algebra isomorphic to the Heisenberg-Weyl Lie algebra, namely
@@
This means that the three dual Lie bialgebras $(\mathfrak{g}^{\ast},\delta^{\ast})$ will have the same dual cocommutator map $\delta^{*}$ arising from the Poincaré Lie algebra (3.5), namely
@@
4 Poisson-Lie deformations and Lie bialgebras
In this Section we study the Poisson-Hopf algebras relevant to the different deformations of the Rikitake systems that we are considering. In particular, for Case A we explicitly compute both of the non-trivial Poisson-Hopf algebras associated to the two non-trivial $(1+1)$-dimensional Poincaré Lie bialgebras previously introduced and show explicitly that in the limit $\eta\to 0$ we recover the Poisson version of (3.5). Afterwards, we study the possible bi-Hamiltonian deformations: we prove that Case B does not admit a bi-Hamiltonian deformation by showing that the Poisson pencil (2.18) does not admit a common cocycle, and finally, we explicitly construct a bi-Hamiltonian deformation for Case AB.
@@ 4.1.1 The ‘book’ group deformation
By using (4.1) we can parametrize a generic element of the group $G^{*}$ in terms of the $(x,y,z)$-coordinates, namely
$\mathcal{G}_{1}=\exp(z\,\varrho(\hat{z}))\exp(y\,\varrho(\hat{y}))\exp(x\,\varrho(\hat{x}))=\begin{pmatrix}e^{-\eta\,z}&0&\eta\,xe^{-\eta z}\\ 0&e^{-\eta\,z}&\eta\,ye^{-\eta z}\\ 0&0&1\end{pmatrix},$ (4.2)
@@
The coproducts for the $(x,y,z)$-coordinates are derived from the group multiplication and read [7, 8]
$\begin{array}[]{l}\Delta_{\eta}(x)=x\otimes e^{\eta\,z}+1\otimes x=x_{1}e^{\eta\,z_{2}}+x_{2},\\ \Delta_{\eta}(y)=y\otimes e^{\eta\,z}+1\otimes y=y_{1}e^{\eta\,z_{2}}+y_{2},\\ \Delta_{\eta}(z)=z\otimes 1+1\otimes z=z_{1}+z_{2},\end{array}$ (4.4)
where we have identified the coordinates of $\mathcal{G}_{1}$ with the tensor space at the left side of the tensor product, and the ones of $\mathcal{G}_{2}$ at the right one.
As the deformed Poisson bracket is always quadratic in terms of the group entries, the application of the compatibility conditions stated by the fact that the deformed coproduct relations (4.4) have to provide a Poisson map leads to the most generic PL structure on the book group, which was explicitly obtained in [27].
From this result, and imposing that the linearisation of such a generic PL structure has to lead to the Poisson version of (3.5), a straightforward computation leads to the unique solution given by the Poisson bracket
@@
4.1.2 The Heisenberg-Weyl deformation